OpenStudy (anonymous):

What is the value of the expression? 2^3 + 5^2

OpenStudy (shinalcantara):

what do you understand of a number raised to a power?

OpenStudy (anonymous):

When having two integers a,b and we have a^b we multiply a*a*a...*a b-times so 2^3=2*2*2=8; 5^2=5*5+25, and their sum would be 25+8=33

OpenStudy (anonymous):

so answer this now 3^3+4^2

OpenStudy (shinalcantara):

33 is the value

OpenStudy (shinalcantara):

it's the same thing. you'll have (3*3*3) + (4*4)